Logged in as "Eric Dufresne"
Mon Jun 19 16:28:12 2006
Time series data from 3h25 am on 6/18 til 6/19 8am
Here are some data taken during 28.5 hours. The first part of the data was taken in 23 hours
without any access to 7ID-C. The ID gap was closed by ED from 13.12 keV to 10.12 keV
shortly before on 6/17/06 at 22h40 (about 5 hours before the start of the time series).
Att. 1-3 show the first 23 hours with 7ID=10.12 keV, mono 10 keV, L5-20 0.5mmx0.5mm.
shows the ring current (red) and diode sum (purple) in 7ID-C at 49.2 m from the source.
shows the beam position X,Y of the 7ID-C X-ray BPM. The beam was quite stable over the
duration of the time serie (25 microns peak to peak). We are not using any feedback.
There is some obvious fast fluctuations in the vertical with a 15 micron peak to peak motion.
This may have been caused by a turbopump and may require further study.
shows a projection of the 7ID front end X-ray BPM P1 and P2 projected on the L5-20 slit
plane, 26.5 m from the source.
Att. 4 and 5 were taken while Don Walko calibrated the monochromator energy. The flux in ion chambers
in 7ID-B and C are displayed now in
shows the beam motion occuring when Don scanned
the energy from ~5 to ~24 keV, i.e. the range of operation of the new monochromator.
We are very happy with the performance of this new monochromator, and it should improve in the near
future as we implement feedback on the second crystal of the monochromator, and learn more to operate